Doug was looking down over Jim's shoulder," That's my new permit number" he said," What's wrong with it"?
"Nothing,Dad,but look" the boy replied,
" If you subtract the square of the last pair of digits from the square of the other three,you get seven less than the number,
And you always say seven's your lucky number?
What was the number.?

Answer by ankor@dixie-net.com(22740)

If you subtract the square of the last pair of digits from the square of the other three, you get seven less than the number,
What was the number.?
:
Let a = the value of the 1st three digits
Let b = the value of the last two digits
:
a^2 - b^2 = 100a+b - 7
:
Could not readily come up with a method to solve this
Wrote a short Basic program and it spit out:
10727
:
The program
10 Print "Subtract the square of the last pair of digits from the square
of the other three, you get seven less than the number":print
20 For a = 100 to 999
30 for b = 10 to 99
40 x = a^2 - b^2
50 y = 100*a + b
60 if x - y = -7 then print a;b
70 next b
80 next a
100 end